# Unlabeled Data or Pre-trained Model: Rethinking Semi-Supervised Learning and Pretrain-Finetuning

### v3

@@ 4.1 Baselines Selection
CIFAR-10/100 78.91 $\pm$ 0.15 70.95 $\pm$ 0.34 77.66 $\pm$ 0.22 80.05 $\pm$ 0.14 79.30 $\pm$ 0.00 87.87 $\pm$ 0.26 92.27 $\pm$ 0.12
ImageNet-30 70.07 $\pm$ 0.26 72.65 $\pm$ 0.70 79.23 $\pm$ 0.29 82.20 $\pm$ 0.38 98.80 $\pm$ 0.00 99.33 $\pm$ 0.12 99.30 $\pm$ 0.00
ImageNet-100 65.11 $\pm$ 0.32 68.43 $\pm$ 0.33 66.85 $\pm$ 0.19 82.81 $\pm$ 0.31 87.70 $\pm$ 0.00 91.37 $\pm$ 0.12 91.70 $\pm$ 0.08
Datasets and Experimental Setup. We evaluate both paradigms on standard open-set benchmarks, including CIFAR-10/100 [27], ImageNet-30 [22], and ImageNet-100 [6]. For CIFAR-10/100, we use 100 labeled images per class from CIFAR-10 as the supervised data and the entire CIFAR-100 dataset as the unsupervised data. For ImageNet-30 and ImageNet-100, we adopt a 50-shot setting, using 50 labeled samples per class from the selected in-distribution (ID) classes. It is important to note that the test set in this setting contains only those classes present in the training data. For the pretrain–finetuning paradigm, the experimental setup and VLM performance remain consistent with those described in the semi-supervised setting. Further details can be found in Appendix A.3.
Experiment Results. As shown in Table 3, zero-shot CLIP and its fine-tuned variants consistently outperform other methods across all datasets, with particularly strong results on ImageNet-30 and ImageNet-100. This advantage stems from their data efficiency—without reliance on noisy unlabeled data, VLMs are less impacted by out-of-distribution (OOD) samples in low-quality datasets. In contrast, SSL methods are more sensitive to such data quality issues, as they require specialized mechanisms to identify and filter out OOD samples. This reliance on additional OOD handling procedures inevitably affects both training efficiency and final performance, often leading to suboptimal results. Moreover, unlike SSL paradigms, pretrain–finetuning approaches benefit from extensive large-scale pretraining followed by targeted fine-tuning, resulting in consistently lower performance variance and a higher ceiling that scales with the amount of labeled data. Based on the above analysis, we conclude that the pretrain–finetuning paradigm is more suitable in scenarios where unlabeled data contains previously unseen classes.
Takeaway 2: When the unlabeled data contains out-of-distribution classes, pre-trained VLMs can be leveraged to avoid unlabeled training, capitalizing on their extensive pre-trained knowledge and rapid adaptation capabilities. Additionally, pre-trained VLMs offer advantages such as robust result stability and high performance.
4.4 SSL vs. VLM Finetuning in Open-World Semi-Supervised Setting
@@
FixMatch [48] 71.5 50.4 49.5 39.6 23.5 20.3 65.8 36.7 34.9
DS3[17] 77.6 45.3 40.2 55.1 23.7 24.0 71.2 32.5 30.8
CGDL [49] 72.3 44.6 39.7 49.3 22.5 23.6 67.3 33.8 31.9
DTC [19] 53.9 39.5 38.3 31.3 22.9 18.3 25.6 20.8 21.3
RankStats [18] 86.6 81.0 82.9 36.4 28.4 23.1 47.3 28.7 40.3
SimCLR [8] 58.3 63.4 51.7 28.6 21.1 22.3 39.5 35.7 36.9
ORCA [7] 88.2 90.4 89.7 66.9 43.0 48.1 89.1 72.1 77.8
NACH [16] 89.5 92.2 91.3 68.7 47.0 52.1 91.0 75.5 79.6
OpenLDN [44] 95.7 95.1 95.4 73.5 46.8 60.1 89.6 68.6 79.1
OpenCon [50] 89.3 91.1 90.4 69.1 47.8 52.7 90.6 80.8 83.8
OwMatch [39] 93.0 95.9 94.4 74.5 55.9 65.1 91.7 72.0 81.8
OwMatch+ [39] 96.5 97.1 96.8 80.1 63.9 71.9 91.5 79.6 85.5
ZSCLIP [43] 88.0 88.1 79.3 57.4 53.1 46.0 89.5 87.9 87.7
CoOp [57] 91.1 90.4 81.3 65.4 51.0 48.4 89.7 83.2 85.5
PromptSRC [25] 93.4 91.1 83.9 73.0 63.2 59.7 91.7 89.5 89.5
@@
Datasets and Experimental Setup. We conduct evaluations on CIFAR-10/100 [27], and ImageNet-100 [6]. In each dataset, the classes are divided into known (base) and new classes, with the proportion of new classes set to 50% by default. A portion of the samples from the known classes are labeled according to a given label ratio, while the remaining part, together with all the new class data, formed the unlabeled set. This setup requires the model to simultaneously distinguish and classify both base and new classes. Within the context of pretrain-finetuning paradigm, this is commonly known as the Base-to-New setting. For more detailed settings, please refer to Appendix A.4.
@@
Experiment Results. In the open-world setting, SSL methods aim to identify unseen classes from unlabeled data during training, typically using clustering-based pseudo-labeling, while evaluating both base and novel classes at test time. As shown in Table 4, these methods achieve the best performance on CIFAR-10 and CIFAR-100, further supporting the conclusions drawn from the semi-supervised setting analysis. In contrast, on the ImageNet-100 dataset, zero-shot CLIP already approaches the performance of OwMatch on base classes and significantly outperforms it on novel classes. This can be attributed to the richer feature representations captured by pre-trained models as image resolution increases, enabling them to better leverage prior knowledge. Notably, pretrain-finetuning approaches exhibit a smaller performance gap between base and novel classes on ImageNet-100 compared to SSL methods, highlighting their strong generalization capabilities, which allow for rapid adaptation to new classes.
@@
Takeaway 3: In the open-world setting, SSL methods perform better on low-resolution datasets, while pre-trained VLMs demonstrate balanced performance on both base and new classes in high-resolution datasets. Combined with our open-set analysis, we conclude that the pretrain-finetuning paradigm is better suited for scenarios where unlabeled data may contain previously unseen classes.
4.5 SSL vs. VLM Finetuning in Long-Tailed Semi-Supervised Setting
@@
Table 5: Test accuracy of SSL and pretrained VLMs under consistent imbalance ratio setting on CIFAR10-LT and CIFAR100-LT datasets.
@@
$N_{1}=500$ $M_{1}=4000$ $N_{1}=1500$ $M_{1}=3000$ $N_{1}=500$ $M_{1}=4000$ $N_{1}=1500$ $M_{1}=3000$ $N_{1}=50$ $M_{1}=400$ $N_{1}=150$ $M_{1}=300$ $N_{1}=50$ $M_{1}=400$ $N_{1}=150$ $M_{1}=300$
Supervised 47.3 $\pm$ 0.95 61.9 $\pm$ 0.41 44.2 $\pm$ 0.33 58.2 $\pm$ 0.29 29.6 $\pm$ 0.57 46.9 $\pm$ 0.22 25.1 $\pm$ 1.14 41.2 $\pm$ 0.15
w/ LA [37] 53.3 $\pm$ 0.44 70.6 $\pm$ 0.21 49.5 $\pm$ 0.40 67.1 $\pm$ 0.78 30.2 $\pm$ 0.44 48.7 $\pm$ 0.89 26.5 $\pm$ 1.31 44.1 $\pm$ 0.42
FixMatch + LA [37] 75.3 $\pm$ 2.45 82.0 $\pm$ 0.36 67.0 $\pm$ 2.49 78.0 $\pm$ 0.91 47.3 $\pm$ 0.58 58.6 $\pm$ 0.36 41.4 $\pm$ 0.93 53.4 $\pm$ 0.32
w/ DARP [26] 76.6 $\pm$ 0.92 80.8 $\pm$ 0.62 68.2 $\pm$ 0.24 76.7 $\pm$ 1.35 50.4 $\pm$ 0.65 59.0 $\pm$ 0.53 43.6 $\pm$ 1.05 55.3 $\pm$ 0.84
w/ CReST+ [51] 76.7 $\pm$ 1.31 81.1 $\pm$ 0.57 70.9 $\pm$ 1.14 77.9 $\pm$ 0.71 44.0 $\pm$ 0.66 57.1 $\pm$ 1.10 40.6 $\pm$ 0.52 52.3 $\pm$ 0.30
w/ DASO [40] 77.9 $\pm$ 0.88 82.5 $\pm$ 0.60 71.0 $\pm$ 1.68 79.0 $\pm$ 2.23 50.1 $\pm$ 0.75 61.0 $\pm$ 0.72 43.5 $\pm$ 0.81 56.1 $\pm$ 0.72
FixMatch + ACR [52] 81.6 $\pm$ 0.19 84.1 $\pm$ 0.20 77.0 $\pm$ 1.19 82.0 $\pm$ 0.41 51.1 $\pm$ 0.80 61.3 $\pm$ 0.26 46.1 $\pm$ 0.34 56.5 $\pm$ 0.33
FixMatch + CPE [35] 80.7 $\pm$ 0.96 84.4 $\pm$ 0.29 76.8 $\pm$ 0.53 82.3 $\pm$ 0.36 50.6 $\pm$ 0.45 59.8 $\pm$ 0.13 43.6 $\pm$ 0.88 55.8 $\pm$ 0.65
FixMatch + CCL [58] 84.5 $\pm$ 0.38 86.2 $\pm$ 0.35 81.5 $\pm$ 0.99 84.0 $\pm$ 0.21 53.5 $\pm$ 0.49 63.5 $\pm$ 0.39 46.8 $\pm$ 0.45 57.5 $\pm$ 0.55
ZSCLIP [43] 79.30 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp [57] 80.73 $\pm$ 1.90 84.20 $\pm$ 0.50 80.60 $\pm$ 1.79 82.23 $\pm$ 0.75 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50 44.83 $\pm$ 0.45 52.57 $\pm$ 0.12
PromptSRC [25] 86.50 $\pm$ 0.37 88.70 $\pm$ 0.08 86.03 $\pm$ 0.48 87.67 $\pm$ 0.42 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34 58.03 $\pm$ 0.46 62.83 $\pm$ 0.48
Method STL10-LT STL10-LT ImageNet-127
$\gamma=10$ $\gamma=20$ $\gamma=\text{None}$
$N_{1}=150$ $M=100k$ $N_{1}=450$ $M=100k$ $N_{1}=150$ $M=100k$ $N_{1}=450$ $M=100k$ 32 $\times$ 32 64 $\times$ 64
   w/ DARP [26] 66.9 $\pm$ 1.66 75.4 $\pm$ 1.05 59.9 $\pm$ 2.17 72.3 $\pm$ 0.60 30.5 42.5
   w/ CReST+ [51] 56.0 $\pm$ 3.13 71.5 $\pm$ 0.96 55.4 $\pm$ 3.03 68.5 $\pm$ 1.15 32.5 44.7
   w/ DASO [40] 61.3 $\pm$ 1.84 78.4 $\pm$ 0.96 68.5 $\pm$ 1.78 75.3 $\pm$ 0.44 40.9 55.9
   w/ CCL [58] 79.1 $\pm$ 0.43 84.8 $\pm$ 1.05 77.1 $\pm$ 0.33 83.1 $\pm$ 0.18 61.5 67.8
ZSCLIP [43] 97.4 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp [57] 97.77 $\pm$ 0.19 98.13 $\pm$ 0.17 97.73 $\pm$ 0.25 97.83 $\pm$ 0.09 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50
PromptSRC [25] 98.27 $\pm$ 0.05 98.47 $\pm$ 0.05 98.13 $\pm$ 0.12 98.30 $\pm$ 0.08 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34
@@
Datasets and Experimental Settings. We compare different methods on four long-tailed datasets: CIFAR-10-LT [27], CIFAR-100-LT [27], STL-10-LT [9], and ImageNet-127 [12]. For synthetic long-tailed datasets (CIFAR, STL-10), the class-wise labeled/unlabeled sample counts follow an exponential decay controlled by imbalance ratios $\gamma_{l}$ and $\gamma_{u}$ [52]. For ImageNet-127, a naturally long-tailed dataset, we follow [58] and use 10% of the training data as labeled. To ensure fair comparison, we adopt the same splits and protocols as SSL methods [39], and report performance of VLMs in a few-shot setting where the number of shots is determined by the smallest labeled class (e.g., 5-shot for CIFAR-10 when $\gamma_{l}=100$ and $N_{1}=500$). Full setup details and shot derivations are available in Appendix A.5.
@@
Experiment Results. As shown in Table 5, in the long-tailed semi-supervised setting, SSL methods are designed to ensure that the learning process effectively captures all classes, even when the distribution is dominated by a few head classes. However, as shown in Table 1, their performance often declines in the same datasets compared to standard SSL, highlighting their ongoing vulnerability to the challenges of long-tailed distributions. In contrast, the evaluation results of pre-trained VLMs underscore their efficiency, which can be fine-tuned with only a small amount of labeled data and still perform well, even on low-resolution datasets, a common challenge for these models. As shown in Table 6, pre-trained models perform well on highly imbalanced datasets like STL-10 but struggle on semantically coarse datasets such as ImageNet-127 [12]. ImageNet-127 is obtained by collapsing the 1,000 fine-grained ImageNet-1K classes into 127 coarse-grained categories according to their positions in the WordNet semantic hierarchy, yielding a long-tailed distribution (e.g., grouping “abacus,” “acoustic guitar,” and “analog clock” under the broader category “device”). Zero-shot CLIP evaluations reveal that this hierarchical coarsening injects semantic ambiguity, making it challenging for pretrained VLMs to confidently associate a fine-grained concept with a specific image. Although downstream fine-tuning can partially mitigate this issue, the underlying semantic constraints remain. Detailed analysis and visualization are shown in Appendix C.
Takeaway 4: The low data requirements of pre-trained models enable them to effectively handle long-tailed distributions, offering a promising approach to addressing the inherent imbalance in many real-world datasets. Meanwhile, SSL methods remain well-suited for tasks characterized by less distinct semantic boundaries between categories, where learning directly from data representations proves advantageous.
5 Discussion
Potential for combining SSL and pre-trained models. Integrating SSL with pre-trained models offers new opportunities for semi-supervised classification. While SSL leverages unlabeled data to reduce labeling costs, pre-trained models excel in zero-shot learning and generalization. Their combination can effectively address challenges such as few-shot learning, class imbalance, and other specialized scenarios. For instance, PromptKD [34] introduces a two-stage, prompt-driven knowledge distillation framework that aligns teacher and student models through learnable prompts, achieving effective knowledge transfer using a diverse mix of ID and OOD unlabeled images, resulting in significant improvements across various datasets. Similarly, UPL [23] automatically generates high-quality prompts from unlabeled data, reducing the dependence on large-scale labeled datasets and improving overall model fit. GRIP [36] represents another promising direction, proposing a pseudo-label-based fine-tuning method that effectively combines SSL principles with VLMs. Clearly, leveraging pseudo-labeled data, semi-supervised learning concepts, and the generalization strengths of VLMs offers a promising path for enhancing performance across a wide range of downstream tasks, prompting a fresh perspective on hybrid learning strategies.

### v4

@@ 4.3 Comparisons in Open-Set SSL Setting
CIFAR-10/100 78.91±0.15 70.95±0.34 77.66±0.22 80.05±0.14 79.30±0.00 87.87±0.26 92.27±0.12
ImageNet-30 70.07±0.26 72.65±0.70 79.23±0.29 82.20±0.38 98.80±0.00 99.33±0.12 99.30±0.00
ImageNet-100 65.11±0.32 68.43±0.33 66.85±0.19 82.81±0.31 87.70±0.00 91.37±0.12 91.70±0.08
Datasets and Experimental Setup. We evaluate both paradigms on standard open-set benchmarks, including CIFAR-10/100 (Krizhevsky et al., 2009), ImageNet-30 (Hendrycks et al., 2019), and ImageNet-100 (Cao et al., 2021). For CIFAR-10/100, we use 100 labeled images per class from CIFAR-10 as the supervised data and the entire CIFAR-100 dataset as the unsupervised data. For ImageNet-30 and ImageNet-100, we adopt a 50-shot setting, using 50 labeled samples per class from the selected in-distribution (ID) classes. It is important to note that the test set in this setting contains only the classes seen during training. For pre-trained models, the experimental setups are consistent with those in the semi-supervised setting. Further details can be found in Appendix B.3.
Experiment Results. As shown in Table 3, zero-shot CLIP and its fine-tuned variants consistently outperform other methods across all datasets, with particularly strong results on ImageNet-30 and ImageNet-100. This advantage stems from their data efficiency—without reliance on noisy unlabeled data, VLMs are less impacted by out-of-distribution (OOD) samples in low-quality datasets. In contrast, SSL methods are more sensitive to such data quality issues, as they require specialized mechanisms to identify and filter out OOD samples. This reliance on additional OOD handling procedures inevitably affects both training efficiency and final performance, often leading to suboptimal results. Moreover, pre-trained models derive distinct advantages from two key stages: first, they undergo extensive large-scale pre-training on diverse datasets, and second, they are refined through targeted fine-tuning on task-specific data. This two-step process enables them to achieve consistently lower performance variance across different data distributions and a higher performance ceiling. Based on the analysis, we conclude that pretrained models are better suited for scenarios where unlabeled data contain previously unseen classes, and that the impact of noisy data is far more severe than that of “adaptation bias”.
Takeaway 2: When the unlabeled data contains out-of-distribution classes, pre-trained models can be leveraged to avoid unlabeled training, capitalizing on their pre-trained knowledge and rapid adaptation capabilities. Additionally, pre-trained models offer advantages such as robust result stability and high performance.
4.4 Comparisons in Open-World SSL Setting
@@
FixMatch 71.5 50.4 49.5 39.6 23.5 20.3 65.8 36.7 34.9
DS3L 77.6 45.3 40.2 55.1 23.7 24.0 71.2 32.5 30.8
CGDL 72.3 44.6 39.7 49.3 22.5 23.6 67.3 33.8 31.9
DTC 53.9 39.5 38.3 31.3 22.9 18.3 25.6 20.8 21.3
RankStats 86.6 81.0 82.9 36.4 28.4 23.1 47.3 28.7 40.3
SimCLR 58.3 63.4 51.7 28.6 21.1 22.3 39.5 35.7 36.9
ORCA 88.2 90.4 89.7 66.9 43.0 48.1 89.1 72.1 77.8
NACH 89.5 92.2 91.3 68.7 47.0 52.1 91.0 75.5 79.6
OpenLDN 95.7 95.1 95.4 73.5 46.8 60.1 89.6 68.6 79.1
OpenCon 89.3 91.1 90.4 69.1 47.8 52.7 90.6 80.8 83.8
OwMatch 93.0 95.9 94.4 74.5 55.9 65.1 91.7 72.0 81.8
OwMatch+ 96.5 97.1 96.8 80.1 63.9 71.9 91.5 79.6 85.5
ZSCLIP 88.0 88.1 79.3 57.4 53.1 46.0 89.5 87.9 87.7
CoOp 91.1 90.4 81.3 65.4 51.0 48.4 89.7 83.2 85.5
PromptSRC 92.9 90.0 83.8 72.3 62.8 58.9 91.2 89.7 89.4
@@
Datasets and Experimental Setup. We conduct evaluations on CIFAR-10/100 (Krizhevsky et al., 2009), and ImageNet-100 (Cao et al., 2021). In each dataset, the classes are divided into known (base) and new classes, with the proportion of new classes set to 50% by default. A portion of the samples from the known classes are labeled according to a given label ratio, while the remaining part, together with all the new class data, formed the unlabeled set. This setup requires the model to distinguish and classify both base and new classes simultaneously. For more detailed settings, please refer to Appendix B.4.
@@
Experiment Results. In the open-world setting, SSL methods aim to identify unseen classes from unlabeled data during training, typically using clustering-based pseudo-labeling, while evaluating both base and novel classes at test time. As shown in Table 4, SSL methods achieve the best performance on CIFAR-10 and CIFAR-100 datasets, indicating that specially tailored losses outperform pretrained models affected by “adaptation bias”, although requiring substantial GPU training time. In contrast, on the ImageNet-100 dataset, zero-shot CLIP already approaches the performance of OwMatch on base classes and significantly outperforms it on novel classes. This can be attributed to the richer feature representations captured by pre-trained models as image resolution increases, thereby reducing “adaptation bias”. Notably, fine-tuned models exhibit a smaller performance gap between base and novel classes on ImageNet-100 compared to SSL methods, highlighting their strong generalization capabilities. In summary, pre-trained models that overcome adaptation bias offer advantages such as rapid adaptability and strong generalization, whereas SSL methods require careful design to effectively handle both novel and unseen classes.
@@
Takeaway 3: In the open-world setting, SSL methods perform better on low-resolution datasets, while pre-trained VLMs demonstrate balanced performance on both base and new classes in high-resolution datasets. Combined with our open-set analysis, we conclude that the pre-trained models are better suited for scenarios where unlabeled data may contain unseen classes.
4.5 Comparisons in Long-Tailed SSL Setting
@@
Table 5: Performance of SSL and VLMs under consistent imbalance ratio setting.
@@
$N_{1}=500$
$M_{1}=4000$
$N_{1}=1500$
$M_{1}=3000$
$N_{1}=500$
$M_{1}=4000$
$N_{1}=1500$
$M_{1}=3000$
$N_{1}=50$
$M_{1}=400$
$N_{1}=150$
$M_{1}=300$
$N_{1}=50$
$M_{1}=400$
$N_{1}=150$
$M_{1}=300$
Supervised 47.3±0.95 61.9±0.41 44.2±0.33 58.2±0.29 29.6±0.57 46.9±0.22 25.1±1.14 41.2±0.15
w/ LA 53.3±0.44 70.6±0.21 49.5±0.40 67.1±0.78 30.2±0.44 48.7±0.89 26.5±1.31 44.1±0.42
FixMatch + LA 75.3±2.45 82.0±0.36 67.0±2.49 78.0±0.91 47.3±0.58 58.6±0.36 41.4±0.93 53.4±0.32
w/ DARP 76.6±0.92 80.8±0.62 68.2±0.24 76.7±1.35 50.4±0.65 59.0±0.53 43.6±1.05 55.3±0.84
w/ DASO 77.±0.88 82.5±0.60 71.0±1.68 79.0±2.23 50.1±0.75 61.0±0.72 43.5±0.81 56.1±0.72
FixMatch + ACR 81.6±0.19 84.1±0.20 77.0±1.19 82.0±0.41 51.1±0.80 61.3±0.26 46.1±0.34 56.5±0.33
FixMatch + CPE 80.7±0.96 84.4±0.29 76.8±0.53 82.3±0.36 50.6±0.45 59.8±0.13 43.6±0.88 55.8±0.65
FixMatch + CCL 84.5±0.38 86.2±0.35 81.5±0.99 84.0±0.21 53.5±0.49 63.5±0.39 46.8±0.45 57.5±0.55
ZSCLIP 79.30±0.00 46.00±0.00
CoOp 80.73±1.90 84.20±0.50 80.60±1.79 82.23±0.75 49.83±1.04 55.23±0.50 44.83±0.45 52.57±0.12
PromptSRC 86.50±0.37 88.70±0.08 86.03±0.48 87.67±0.42 60.33±0.17 64.37±0.34 58.03±0.46 62.83±0.48
@@
Datasets and Experimental Setup. We compare different methods on four long-tailed datasets: CIFAR-10-LT (Krizhevsky et al., 2009), CIFAR-100-LT (Krizhevsky et al., 2009), STL-10-LT (Coates et al., 2011), and ImageNet-127 (Fan et al., 2022). For synthetic long-tailed datasets (CIFAR, STL-10), the class-wise labeled/unlabeled sample counts follow an exponential decay controlled by imbalance ratios $\gamma_{l}$ and $\gamma_{u}$ (Wei & Gan, 2023). For ImageNet-127, a naturally long-tailed dataset, we follow (Zhou et al., 2024) and use 10% of the training data as labeled. To ensure fair comparison, we adopt the same splits and protocols as SSL methods (Niu et al., 2024), and report performance of pre-trained models in a few-shot setting where the number of shots is determined by the smallest labeled class (e.g., 5-shot for CIFAR-10 when $\gamma_{l}=100$ and $N_{1}=500$). Full setup details and shot derivations are available in Appendix B.5.
@@
Experiment Results. As shown in Table 5, in the long-tailed semi-supervised setting, SSL methods are designed to ensure that the learning process effectively captures all classes, even when the distribution is dominated by a few head classes. However, as shown in Table 1, their performance often declines in the same datasets compared to standard SSL, highlighting their ongoing vulnerability to the challenges of long-tailed distributions. In contrast, the evaluation results of pre-trained models underscore their efficiency, which can be fine-tuned with only a small amount of labeled data and still perform well, even on low-resolution datasets, a common challenge for these models.
We also conducted experiments about STL-10 and Imagenet-127 datasets (detailed results are shown in Appendix C), and concluded that pre-trained models perform well on imbalanced datasets but struggle with semantically coarse tasks due to ambiguous class definitions. Although downstream fine-tuning can partially mitigate this issue, the underlying semantic constraints remain. We provide a more in-depth explanation and visualization of this phenomenon in Appendix C. Ultimately, we concluded that the highly abstract nature of downstream task categories inherently leads to “adaptation bias”, a fundamental issue that negatively impacts the performance of VLMs.
Takeaway 4: The low data requirements of pre-trained models allow them to effectively address long-tailed distributions, offering a promising solution to real-world data imbalance. However, when the semantic boundaries of the downstream tasks are not clear, pre-trained VLMs’ adaptability will decline.
4.6 Comparisons between VLMs and MLLMs
Datasets and Experimental Setup. We compare the performance of CLIP with existing MLLM approaches such as Qwen and GPT-4o-mini on the CIFAR-10, STL-10, ImageNet-30, and ImageNet-100 datasets. The detailed prompts for MLLMs are provided in Appendix D.
Experiment Results. Intuitively, increasing the number of model parametercan improve classification accuracy; however, this phenomenon is primarily observed in CLIP models. For MLLMs, which are generally assumed to handle image classification tasks with ease, often underperform compared to CLIP, and simply increasing the model size can even lead to performance degradation (see Table 6). In contrast, GPT-4o-min demonstrates strong performance on these datasets, suggesting that its developers have explicitly enhanced the visual capability of its encoder. Regarding the observation that larger-parameter MLLMs do not consistently outperform CLIP-based models in image classification tasks, our case studies in Appendix D reveal that MLLMs often fail to identify which subjects in an image to focus on. This can be attributed to their autoregressive training paradigm, which is less conducive to effective vision-language alignment. Taken together, our analysis indicates that while scaling up parameters enhances a model’s training capacity, it does not necessarily improve classification ability—a factor that has been overlooked in prior Qwen research.
Takeaway 5: Although large language models possess stronger semantic understanding, their advantages in purely visual classification tasks remain limited. Therefore, the ability to effectively extract and leverage visual features should be emphasized as a key direction for future research, not only for VLMs like CLIP but also for MLLMs such as the Qwen series.
Table 6: Classification performance comparison between VLMs and MLLMs.
Method CIFAR-10 ImageNet-30 STL-10 ImageNet-100
CLIP-ViT-B/16 79.30 98.40 97.40 87.70
CLIP-ViT-L/14 86.80 99.31 98.90 89.70
Qwen2.5-VL-7B-Instruct 74.51 97.63 92.27 65.92
Qwen2.5-VL-32B-Instruct 67.06 99.00 87.52 71.60
GPT-4o-mini 89.82 99.43 98.31 85.36
5 Conclusions and Discussions
Conclusions We conduct a comprehensive and fair comparison of the strengths and weaknesses of SSL methods and pre-trained models. The key insights are as follows:
  1. 1.
    SSL methods suffer from high training cost overhead and usually suffer lower performance (Takeaway 1). Moreover, they suffer from severe performance degradation in open environments, such as class space and distribution changes between labeled and unlabeled data, whereas pre-trained models generalize well owing to their pre-trained knowledge (Takeaway 2, 3).
  2. 2.
    Though fine-tuning pre-trained models typically enable rapid performance gains, they may suffer from “adaptation bias” when downstream task data differs from pre-trained data distribution (Takeaway 1, 4). For example, the CLIP model, which is pre-trained on data with $224\times 224$ resolutions, achieves poor performance for low-resolution data, such as CIFAR-10.
Table 7: Performance of representative methods in the fusion paradigm on the Flowers102 dataset.
Method Domain Data SSL Open-World
CLIP Radford et al. (2021) Zero-shot 72.08 77.80
FPL (Menghini et al., 2023) Few-shot + Unlabeled 75.96 80.97
IFPL (Menghini et al., 2023) 78.68 82.08
GRIP (Menghini et al., 2023) 83.60 86.26
PromptKD (Li et al., 2024) 98.23 86.27
CPL (Zhang et al., 2024) 89.66 87.35
Discussion 1: Potential for combining SSL and pre-trained models. Integrating SSL with pre-trained models presents a promising direction for enhancing classification performance, combining the ability to exploit unlabeled data with strong generalization and zero-shot capabilities. This synergistic effect will help tackle challenges such as few-shot learning, class imbalance, and other domain-specific scenarios. For example, GRIP (Menghini et al., 2023) and CPL (Zhang et al., 2024) bridge SSL methods and VLMs through a pseudo-label-based fine-tuning strategy. Futhermore, PromptKD (Li et al., 2024) introduces a two-stage framework that distills knowledge from teacher to student via learnable prompts and unlabeled data. As shown in Table 7, incorporating unlabeled data and using labels generated by pre-trained models not only leverages the rich prior knowledge embedded in pre-trained models, but also enables semi-supervised learning algorithms to effectively mine domain-specific knowledge from unlabeled data, thereby achieving substantial improvements in model performance and generalization.